# Positive Effects of Financing

### v2

@@ 4.1.1 Objective Function for the Stochastic Model
For the calibration we consider the expectation operator
@@ 4.1.2 Abatement Policy - Classical DICE Model
The model calibration determines the optimal abatement policy $t\mapsto\mu(t)$ and optimal savings rate $s(t)$ that maximizes the integrated discounted welfare (1), (2).
In the time-discrete classical DICE model the calibration determines the optimal parameters $i\mapsto\left(\mu(t_{i}),s(t)\right)$. For a model time horizon of $T=500\leavevmode\nobreak\ \text{years}$ and a time-discretization step of $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$ this results in 499 free parameters for the abatement policy (and another corresponding set for the savings rate).
@@ 4.2 Cost
To characterize the resulting optimal mitigation paths, we utilise derived quantities. A prominent example of such a derived quantity is the social cost of carbon, $SCC(t)$. The $SCC$ is the time-$t$ marginal cost of emitting one additional unit of carbon. Its unit is $[SCC]=\frac{\mathrm{USD}}{\mathrm{tCO}_{2}}$.
@@ 4.2.1 Distribution of Cost
We analyse the temporal distribution of the cost $C(t)$, which is the sum of the abatement cost $C_{\mathrm{A}}$ and the damage cost $C_{\mathrm{D}}$
@@
To gain a better understanding, we decompose how changes of the abatement policy $\mu$ translate into changes of the cost structure and how these enter into the calibration’s objective function.
@@ 4.2.2 Sensitivity of Damage and Abatement Cost to Policy Changes
The optimal abatement policy $\mu$, representing the equilibrium state of the model, fulfils
@@
From eq. 6 we see why the model calibration may result in some intergenerational inequality: Let $s$ be a fixed time at which a change in the abatement policy $\mu(s)$ is considered. Since $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}\mu(s)>0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}\mu(s)<0$ we see from eq. 6 that the model equates the increment of the cost of damage to the increment of the cost of abatement, both weighted by the value-to-cost-sensitivity $\mathrm{d}V(t)/\mathrm{d}C(t)\cdot N(0)/N(t)$. We have
@@
A change in the abatement policy in time $s$ affects the abatement cost $C_{\mathrm{A}}(s)$ only at time $s$, whereas it affects damage cost $C_{\mathrm{D}}(t)$ for possibly all $t\geq s$. This motivates to consider the damage cost per abatement cost induced by a change in the abatement policy $\mu(s)$, i.e., $\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\big{/}\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}$. As the model calibration is weighting cost by the cost-to-value weight, we consider the corresponding weighted damage-cost per abatement-cost sensitivity. The minus in (8) just accounts for the fact that abatement cost sensitivity and damage cost sensitivity have opposite signs.
@@
From Equation 5 we see that for fixed $s$ the integral over $t\mapsto\rho(s;t)$ equals $1$, i.e., $\rho_{\mathrm{D}}(s)$ can be interpreted as a density. Hence, we may interpret $\int_{0}^{\infty}t\cdot\rho_{\mathrm{D}}(s;t)\ \mathrm{d}t$ as the expected time of damage cost per avoided abatement cost in time $s$.
@@
In the classical DICE 2016 model the expected time of damage cost per avoided abatement cost is approximately 70 years after the time of abatement (Figure 8), indicating that balancing marginal abatement and damage cost is a matter of intergenerational equity.
@@
4.2.3 Cost per GDP
To allow for a netting inside a generation one may consider the (discounted) average value over a generation lifetime. For a given time $s$ let $T^{\mathrm{L}}(s)$ denote the expected lifetime of a generation. We then consider the lifetime average cost and GDP.
@@ 4.3.1 Stochastic Interest Rates
The classical way in which interest rates are modelled in an IAM is via a discount factor. A time-$t$ value $V(t)$ is derived in the model, then discounted and aggregated to a final value forming the objective function
$\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\mathrm{d}t\text{.}$
We model stochastic interest rates $r$. Our implementation allows to use a general discrete forward rate model (LIBOR Market Model). Our results were conducted with a classical Hull-White model. The model provides the stochastic numeráire $N$ (and thus discount factor) as well as the stochastic forward rate $FR$. Setting the volatility of the interest rate to zero recovers the deterministic model with a possibly time-dependent interest rate (term-structure of interest rates). For exemplary parameterization, we set the mean reversion speed to $a(t)=0.02$ and the volatility to $\sigma(t)=0.4\%$. While our result should be interpreted as a qualitative result, showing that stochastic interest rates generate adverse effects, the interest rate scenarios generated by our parameter choice are realistic, rather conservative: the 1st and 99th percentile of the generated interest rates at t = 100 years are -1 % and 5 %, respectively. For a corresponding visualization, see [39].
@@
$\begin{split}\mathrm{E}\left(\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\mathrm{d}t\right)&\ =\ \int_{0}^{T}V(t)\mathrm{E}\left(\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\right)\mathrm{d}t\\ &\ =\ \int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\bar{r}(s)\mathrm{d}s\right)\mathrm{d}t\text{,}\end{split}$ (10)
with $\bar{r}(t)\ =\ -\frac{\partial}{\partial t}\log\left(\mathrm{E}\left(\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\right)\right)$.
Due to the lack of a feedback that allows to adjust the abatement path and resulting damages, adding stochastic interest rates does not yet introduce changes in the model dynamics. This changes however, if one assumes that the abatement policy is adapted to the stochastic changes, here the change in interest rate level, i.e., if one allows $\mu$ to be an (adapted) stochastic process. [^2] Our choice to make interest rates stochastic is only exemplary. The interest rate is an important factor in linking present abatement cost to the avoided future damage cost. While one may introduce stochasticity in many other state variables, the use of an abatement policy adapted to the economic factors (interest rates) will already lead to all other state variables becoming stochastic.
We use a classical Hull-White model, that is
$\begin{split}\mathrm{d}r(t)&\ =\ (\theta(t)-a(t)r(t))\ \mathrm{d}t\ +\ \sigma(t)\ \mathrm{d}W(t)\text{,}\quad r(t_{0})\ :=\ r_{0}\text{,}\\ \mathrm{d}N(t)&\ =\ r(t)\ \mathrm{d}t\text{,}\end{split}$ (11)
where an exact time-discretization scheme [47] was used. The model’s mean-reversion level $\theta$ is such that the numéraire-relative zero coupon bond prices are martingales and their expectation agrees with the classical non-stochastic model. In other words, our stochastic interest rate model is such that the optimal deterministic abatement strategy would result in the same optimal emission pathway as the model with deterministic interest rates.
@@ 4.3.2 Stochastic Abatement Policy
Assuming that a change in the interest rates triggers a re-calibration of the model, a stochastic interest rate implies a stochastic abatement policy, i.e., $\mu$ will become stochastic. The abatement policy $t\mapsto\mu(t)$ becomes a stochastic process that adapts to the changes in interest rates. This reflects the possibility that the abatement policy can be adjusted to the interest rate scenarios.
@@
An approach to determine the optimal exercise strategy is to parameterize $\mu(t)$ as a functional form of $\mathcal{F}_{t}$-measurable random variables and apply a global optimization. As we utilize a one-factor Markovian model for the short rate $r$, we consider a functional form where $\mu$ depends on $t$ and the current interest rate level $r(t)$.
A simple example of a stochastic abatement model is a parametric one, where the abatement speed is a (linear) function of the interest rate level $r$, e.g.,
$\mu(t,\omega)=\min\left(\mu_{0}+\left(a_{0}+a_{1}\cdot r(t,\omega)\right)\cdot t,1.0\right)\text{.}$ (12)
The simplicity of the models allows us to interpret the parameters. We have that $a_{0}$ is the abatement speed at the zero interest rate level and $a_{1}$ describes the change of the abatement speed by interest rate change,
@@
Obviously, the stochastic abatement models may achieve a larger-or-equal welfare or lower-or-equal welfare-risk than the corresponding deterministic model (4) simply due to the additional degree of freedom $a_{1}$ in the abatement optimisation. The calibration of the optimal model (12) results in a negative value for $a_{1}$, reflecting that higher interest rate correspond to slower abatement speed.
@@ 4.3.3 Funding of Abatement Costs
We introduce the option of funding of abatement cost. As abatement of emissions is a planned process of societal relevance, it is reasonable to assume that abatement cost are covered by a loan for which interest rate corresponds to the current discount rate. Our model allows to apply a funding spread, i. e. higher interest rates for loans than for discounting, but this is not considered in this general introduction of the phenomenon. In future work, the concept of the climate beta could be used to introduce an interest rate spread for abatement investments, as introduced in DICE-2023 [38].
@@
In a standard model for risk-neutral valuation, this change would have no effect, as the accruing is compensated by the discounting. However, the change might introduce an effect in the DICE model, due to the way how abatement and damage cost are associated and due to the time-preference included in the utility function.
@@ 4.3.4 Non-Linear Financing Costs
For an unsecured financial cash-flow its present value is defined by a discount factor times the cash-flow. If the future cash-flow is subject to default, the discount factor is lower, reflecting the additional value reduction due to the risk of (partial) default. As default is not an option for future damage cost, it appears as if the risk-free discount factor should apply. However, since no hedging strategy exists, a risk free funding is not possible. Thus additional cost may occur to secure the unsecured funding [41]. Since damage cost may become very large - much larger than funds provided by standard financial markets - it is reasonable to assume that these additional funding cost become (over-proportionally) large for larger cost. We model this by optionally adding non-linear financing cost using a non-linear funding model [41].
This model allows, that the discount factor may depend on the magnitude of the cash-flow. Thus the funding of larger cash-flows requires a premium to compensate for a larger default risk or other frictions. For our application, this means that larger cost get a larger weight.
@@ Acknowledgements
We thank Max Singhoff for exploring the calibration of the DICE model with the ADAM solver used in our numerical experiments.

### v3

@@ 4.2.1 Objective Function for the Stochastic Model
For the optimization we consider the expectation operator
@@ 4.2.2 Abatement Policy - Classical DICE Model
The model optimization determines the optimal abatement policy $\mu(t)$ and optimal savings rate $s(t)$ that maximizes the integrated discounted welfare (2), (3).
In the time-discrete classical DICE model the optimization determines the optimal parameters $i\mapsto\left(\mu(t_{i}),s(t_{i})\right)$. For a model time horizon of $T=500\leavevmode\nobreak\ \text{years}$ and a time-discretization step of $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$ this results in 499 free parameters for the abatement policy (and another corresponding set for the savings rate).
@@ 4.3 Cost
To characterize the resulting optimal mitigation paths, we utilize derived quantities. A prominent example of such a derived quantity is the social cost of carbon, $SCC(t)$. The $SCC$ is the time-$t$ marginal cost of emitting one additional unit of carbon. Its unit is $[SCC]=\frac{\mathrm{USD}}{\mathrm{tCO}_{2}}$.
@@ 4.3.1 Distribution of Cost
We analyze the temporal distribution of the cost $C(t)$, which is the sum of the abatement cost $C_{\mathrm{A}}$ and the damage cost $C_{\mathrm{D}}$
@@
To gain a better understanding, we decompose how changes of the abatement policy $\mu$ translate into changes of the cost structure and how these enter into the optimization’s objective function.
@@ 4.3.2 Sensitivity of Damage and Abatement Cost to Policy Changes
The optimal abatement policy $\mu$, representing the equilibrium state of the model, fulfills
@@
From eq. 7 we see why the model optimization may result in some intergenerational inequality: Let $s$ be a fixed time at which a change in the abatement policy $\mu(s)$ is considered. Since $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}\mu(s)>0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}\mu(s)<0$ we see from eq. 7 that the model equates the increment of the cost of damage to the increment of the cost of abatement, both weighted by the value-to-cost-sensitivity $\mathrm{d}V(t)/\mathrm{d}C(t)\cdot N(0)/N(t)$. We have
@@
A change in the abatement policy in time $s$ affects the abatement cost $C_{\mathrm{A}}(s)$ only at time $s$, whereas it affects damage cost $C_{\mathrm{D}}(t)$ for possibly all $t\geq s$. This motivates to consider the damage cost per abatement cost induced by a change in the abatement policy $\mu(s)$, i.e., $\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\big{/}\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}$. As the model optimization is weighting cost by the cost-to-value weight, we consider the corresponding weighted damage-cost per abatement-cost sensitivity. The minus in (9) just accounts for the fact that abatement cost sensitivity and damage cost sensitivity have opposite signs.
@@
From Equation 6 we see that for fixed $s$ the integral over $t\mapsto\rho(s;t)$ equals $1$, i.e., $\rho_{\mathrm{D}}(s)$ can be interpreted as a density (Figure 8). Hence, we may interpret $\int_{0}^{\infty}t\cdot\rho_{\mathrm{D}}(s;t)\ \mathrm{d}t$ as the expected time of damage cost per avoided abatement cost in time $s$.
@@
In the classical DICE 2016 model the expected time of damage cost per avoided abatement cost is approximately 70 years after the time of abatement. This can be observed in Figure 2, indicating that balancing marginal abatement and damage cost is a matter of intergenerational equity.
@@ 4.4.1 Stochastic Interest Rates
The classical way in which time-preference rates are modeled in an IAM is via a discount factor. A time-$t$ value $V(t)$ is derived in the model, then discounted and aggregated to a final value forming the objective function
$\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\mathrm{d}t\text{.}$
We model stochastic interest rates (time-preference rates) $rho$. Our implementation allows to use a general discrete forward rate model (LIBOR Market Model). Our results were conducted with a classical Hull-White model. The model provides the stochastic numeráire $N$ (and thus discount factor) as well as the stochastic forward rate $FR$. Setting the volatility of the interest rate to zero recovers the deterministic model with a possibly time-dependent interest rate (term-structure of interest rates). For exemplary parametrization, we set the mean reversion speed to $a(t)=0.02$ and the volatility to $\sigma(t)=0.4\%$. While our result should be interpreted as a qualitative result, showing that stochastic interest rates generate adverse effects, the interest rate scenarios generated by our parameter choice are realistic, rather conservative: the 1st and 99th percentile of the generated interest rates at t = 100 years are -1 % and 5 %, respectively. For a corresponding visualization, see [40].
@@
$\begin{split}\mathrm{E}\left(\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\mathrm{d}t\right)&\ =\ \int_{0}^{T}V(t)\mathrm{E}\left(\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\right)\mathrm{d}t\\ &\ =\ \int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\bar{r}(s)\mathrm{d}s\right)\mathrm{d}t\text{,}\end{split}$ (11)
with $\bar{r}(t)\ =\ -\frac{\partial}{\partial t}\log\left(\mathrm{E}\left(\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\right)\right)$.
Due to the lack of a feedback that allows to adjust the abatement path and resulting damages, adding stochastic interest rates does not yet introduce changes in the model dynamics. This changes however, if one assumes that the abatement policy is adapted to the stochastic changes, here the change in interest rate level, i.e., if one allows $\mu$ to be an (adapted) stochastic process. [^2] Our choice to make interest rates stochastic through uncertain time preference rates is only exemplary. The interest rate is an important factor in linking present abatement cost to the avoided future damage cost. While one may introduce stochasticity in many other state variables, the use of an abatement policy adapted to the economic factors (interest rates) will already lead to all other state variables becoming stochastic.
To model the uncertainty of the rate of time-preference $\rho(t)$, we use a classical Hull-White model [39, 55], that is
$\begin{split}\mathrm{d}\rho(t)&\ =\ (\theta(t)-a(t)\rho(t))\ \mathrm{d}t\ +\ \sigma(t)\ \mathrm{d}W(t)\text{,}\quad\rho(t_{0})\ :=\ r_{0}\text{,}\\ \mathrm{d}N(t)&\ =\ \rho(t)\ \mathrm{d}t\text{,}\end{split}$ (12)
where an exact time-discretization scheme [56] was used. Our stochastic interest rate model is parameterized such that the optimal deterministic abatement strategy would result in the same optimal emission pathway as the model with deterministic interest rates.
@@ 4.4.2 Stochastic Abatement Policy
Assuming that a change in the interest rates triggers a re-calculation of the model optimal abatement policy, a stochastic interest rate implies a stochastic abatement policy, i.e., $\mu$ will become stochastic. The abatement policy $t\mapsto\mu(t)$ becomes a stochastic process that adapts to the changes in interest rates. This reflects the possibility that the abatement policy can be adjusted to the interest rate scenarios.
@@
An approach to determine the optimal exercise strategy is to parameterize $\mu(t)$ as a functional form of $\mathcal{F}_{t}$-measurable random variables and apply a global optimization. As we utilize a one-factor Markovian model for the short rate $r$, we consider a functional form where $\mu$ depends on $t$ and the current interest rate level $\rho(t)$.
To obtain a robust optimization, reducing the risk of over-fitting or numerical instabilities, we propose a parametric stochastic abatement model. Here, we use the model
$\mu(t,\omega)=\min\left(\mu_{0}+\left(a_{0}+a_{1}\cdot\rho(t,\omega)\right)\cdot t,1.0\right)\text{,}$ (13)
where the abatement speed is a (linear) function of the interest rate level $\rho$.
The model can be seen as a first-order approximation of general policies, where the abatement is allowed to depend in the level of the discount rate $\rho$.
The simplicity of the model allows us to interpret the parameters. We have that $a_{0}$ is the abatement speed at the zero interest rate level and $a_{1}$ describes the change of the abatement speed by interest rate change,
@@
Obviously, the stochastic abatement models may achieve a larger-or-equal welfare or lower-or-equal welfare-risk than the corresponding deterministic model (5) simply due to the additional degree of freedom $a_{1}$ in the abatement optimization. The optimization of the model (13) results in a negative value for $a_{1}$, reflecting that higher interest rate correspond to slower abatement speed.
@@ 4.4.3 Funding of Abatement Costs
We introduce the option of funding of abatement cost. As abatement of emissions is a planned process of societal relevance, it is reasonable to assume that abatement cost are covered by a loan for which interest corresponds to the current discount rate. Our model allows to apply a funding spread, i. e. higher interest rates for loans than for discounting, but this is not considered in this general introduction of the phenomenon. In future work, the concept of the climate beta could be used to introduce an interest rate spread for abatement investments, comparable to the modifications in DICE-2023 [38].
@@
In a standard risk-neutral valuation, this change would have no effect, as the accruing is compensated by the discounting. However, the change might introduce an effect in the DICE model, due to the way how abatement and damage cost are associated as well as due to the time-preference included in the utility function.
@@ 4.4.4 Non-Linear Financing Costs
For an unsecured financial cash-flow its present value is defined by a discount factor times the cash-flow. If the future cash-flow is subject to default, the discount factor is lower, reflecting the additional value reduction due to the risk of (partial) default. As default is not an option for future damage cost, it appears as if the risk-free discount factor should apply. However, since no hedging strategy exists, a risk free funding is not possible. Thus additional cost may occur to secure the unsecured funding [42]. Since damage cost may become very large - much larger than funds provided by standard financial markets - it is reasonable to assume that these additional funding cost become (over-proportionally) large for larger damages. We model this by optionally adding non-linear financing cost using a non-linear funding model [42].
This model allows, that the discount factor may depend on the magnitude of the cash-flow. Thus the funding of larger cash-flows requires a premium to compensate for a larger default risk or other frictions. For our application, this means that larger damages get a larger weight.
@@
4.4.5 Cost per GDP
To allow for a netting within a generation one may consider the (discounted) average value over a generation lifetime. For a given time $s$ let $T^{\mathrm{L}}(s)$ denote the expected lifetime of a generation. We then consider the lifetime average cost and GDP.
4.5 Convexity
Our numerical analysis shows that a continuous adaptation of the abatement strategy may increase the risk and expectation of intergenerational inequality measures as cost-per-GDP. The expected emissions/damage/cost are higher than the emissions/damage/cost calculated for the expected interest rate level.
Such an effect may come as a surprise. However, it is well known from mathematical finance, known under the name convexity. The effect corresponds to Jensen’s inequality, that the expectation of a convex function of a random variable is higher than the convex function applies to the expectation of the random variable.
Here, the random variable is the exogenous given discount rate, or stochastic process of the interest rates, and the integrated assessment model acts as a convex function. In other words: DICE has convexity in the interest rates once the abatement strategy depends on the interest rate level.
@@ Acknowledgements
We thank Max Singhoff for exploring the optimization of the DICE model with the ADAM solver used in our numerical experiments.
@@
Variation of the Interest Rate Volatility
Supplementary Figure 1: Stochastic DICE Model with Different Interest Rate Volatilities ($\sigma=0.2\%$ (red), $\sigma=0.3\%$ (blue), $\sigma=0.4\%$ (green)). Allowing adaptation of the abatement policy to stochastic interest rates increases intergenerational inequality and exhibits a significant risk of even higher intergenerational inequality (a: emissions, b: abatement, c: damages, and d: total GDP relative cost). Shaded area shows the 10th to 90th percentiles.
Variation of the Interest Rate Level